The amino acid glycine has the following mass percents: 32.00% carbon, 7.14 % hydrogen, 42.63 % oxygen, and 18.66% nitrogen. Determine the empirical formula for glycine.
1
Expert's answer
2020-12-06T13:55:04-0500

Let's say there are 100 grams of the substance. Then there are 32.00 g carbon, 7.14 g hydrogen, 42.63 g oxygen, and 18.66 g nitrogen.

Calculate the number of moles of the atoms.

n(C)= 32/12= 2,67 mole, n(H)=7,14/1=7,14 mole, n(O)=42,63/16=2,66 mole, n(N)=18,66/14=1,33 mole. Moles are related as atoms in a substance. Divide the amounts of substances by the smallest value of the number of moles to get integer values. Receive: С=2, H=5, O=2, N=1 => C2H5NO2.

Answer: empirical formula for glycine is C2H5NO2.